Sir/Madam, can I please get the answer for the following questions:
1.) (x-1)^4+(x-5)^4=82
      Solve for x.
2.) (5+2√6)^(x^2-3)+(5-2√6)^(x^2-3)=10
      Solve for x.

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
Please only one question in one thread.
 
1.)
 
It's easy to check that x = 4, 2 are two solutions to (x - 1)^4 + (x - 5)^4 = 82. 
Now, we can easily find the other values of x. 
(x - 1)^4 + (x - 5)^4 = 82 
==> (x^4 - 4x^3 + 6x^2 - 4x + 1) + (x^4 - 20x^3 + 150x^2 - 500x + 625) = 82 
==> 2x^4 - 24x^3 + 156x^2 - 504x + 544 = 0 
==> x^4 - 12x^3 + 78x^2 - 252x + 272 = 0. 

By long/synthetic division, 
x^4 - 12x^3 + 78x^2 - 252x + 272 
= (x - 4) (x^3 - 8x^2 + 46x - 68) 
= (x - 4)(x - 2)(x^2 - 6x + 34). 

So, we have 
(x - 4)(x - 2)(x^2 - 6x + 34) = 0 
==> x = 4, 2, or x = 3 ± 5i